OpenStudy (anonymous):

1. A ten kilogram rock is in geosynchronous orbit about the earth at the orbital radius of 42 thousand kilometers. calculate the centripetal force

OpenStudy (anonymous):

@agent0smith please help me.. I am comparing the answers with my friend and she got different answers than me.. i dont know why

OpenStudy (anonymous):

I did this, T = 24hrs W= 2 (pi) / T Fc = mrw^2 for w= 2 (pi) / (86400) = (7.27 x 10^-5 then f = (10 kg) (42000000) (7.27 x 10^-5) f = 2.22 N

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

How do you know T = 24?

OpenStudy (loser66):

to me, I just apply F = G m1m2/ r^2

OpenStudy (anonymous):

geosynchronous orbit doesnt that equals 24 hrs which then i convert to seconds?

OpenStudy (loser66):

to my book, it says "A gyosynchoronous satellite is one that stays above the same point on the Earth"

OpenStudy (anonymous):

its not a satellite its an orbit meaning time, i googled and told me it gyosynchoronous orbit is 24hrs

OpenStudy (loser66):

yea, you are right about it but I don't see the link between it with the force. My prof indicated about it when we have to find out the distance between 2 objects to find r with \(r^3= \dfrac{GM T^2}{4\pi^2}\) not to find out the force. I think you overthought it, just F = Gm1m2/r^2 is ok, don't forget r = radius of the earth + orbital raidius

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

@***[ISURU]***